# AgentClinic: a multimodal agent benchmark to evaluate AI in simulated clinical environments

### v3

@@ Results
For bias evaluations we test the most accurate model from the AgentClinic-MedQA framework, GPT-4, as well as Mixtral-8x7B. The normalized accuracy for these experiments are shown in Figure 3 represented as $\text{Accuracy}_{\text{bias}}$ / $\text{Accuracy}_{\text{No Bias}}$ (between 0-100%). GPT-4 and Mixtral-8x7B have an unbiased accuracy equal to $52\%$ and $37\%$ respectively. For GPT-4, we find that cognitive bias results in a larger reduction in accuracy with a normalized accuracy of 92% (absolute accuracy drops from 52% accuracy to 48%) for patient cognitive biases and 96.7% for doctor cognitive biases (absolute drops from 52% to 50.3%). For implicit biases, we find that the patient agent was less affected with a normalized accuracy of 98.6% (absolute drops from 52% to 51.3%), however, the doctor agent was affected as much as cognitive biases with an average of 97.1% (absolute drops from 52% to 50.5%). For cognitive bias, the demonstration was occasionally quite clear in the dialogue, with the patient agent overly focusing on a particular ailment or some unimportant fact. Similarly, the doctor agent would occasionally focus on irrelevant information. However, we find that the implicit bias dialogue does not actually demonstrate observable bias despite having a similar reduction in accuracy for the doctor agent.
@@
Previous work studying cognitive bias in LLMs has shown that GPT-4 is relatively robust to bias compared with other language models[17]. Results from evaluating GPT-4 on AgentClinic-MedQA show only small drops in accuracy with the introduced biases (maximum absolute accuracy reduction of 4%, average reduction of 1.5%). While this reduction can be quite large in the field of medicine, it is a much smaller drop than was observed in previous work (10.2% maximum reduction on BiasMedQA dataset[17]). This might be due to the model being superficially overly-aligned to human values, plausibly leading GPT-4 to not serve as a good model for representing human bias in agent benchmarks as the model may reject to execute on bias instructions (which does not mean that GPT-4 is free of said biases). For example, in our evaluations with gender bias we observed 13 occurrences (out of 107 dialogues) where GPT-4 verbosely rejected to follow through with a bias-related instruction. Mixtral-8x7B saw much larger drops in accuracy than GPT-4 in the presence of bias, and thus might serve as a better model for studying bias.
Bias and patient agent perception
While diagnostic accuracy with GPT-4 did not reduce as much as Mixtral-8x7B, it is also worth investigating the perceived quality of care from the perspective of the patient agent. After the patient-doctor dialogue is completed, we ask every patient agent three questions:
@@
Such patient-agent-centric follow-up queries offer a more fine-grained and multi-faceted characterization of the clinical skills of a language agent—as opposed to static multiple choice benchmarks. The corresponding results are shown in Figure 3. While diagnostic accuracy demonstrates a relatively small drop in accuracy, the patient agent follow-up perceptions tell a different story. Broadly, we find that most patient cognitive biases did not have a strong effect on any of the patient perceptions when compared to an unbiased patient agent except for in the case of self-diagnosis, which had sizeable drops in confidence (4.7 points) and consultation (2 points), and a minor drop in compliance (1 point). However, implicit biases had a profound effect on on all three categories of patient perception, with education bias consistently reducing patient perception across all three categories.
@@
We found that between the implicit biases, sexual orientation bias [^1] Discriminatory bias based on a person’s sexual orientation. had the lowest effect on patient perceptions, followed by racial bias and gender bias. For patient confidence, gender bias is followed by religion socioeconomic, cultural, and education, whereas patient compliance and patient consultation, it is followed by cultural, socioeconomic, religion, and education. While it is not quantifiable, we decided to ask two biased patient agents who provided low rating with education and gender biases for compliance why they provided low ratings. These patient agents had the same symptoms and diagnosis and only differed in bias presentation.
Figure 6: Comparison of accuracy of models on MedQA and AgentClinic-MedQA. We find that MedQA accuracy is not predictive of accuracy on AgentClinic-MedQA due to the additional complexity of dialogue.
With an education bias the patient agent responds: "I would rate my likelihood to consult again with this doctor as a two because, despite their diagnosis, I felt uneasy about their qualifications due to the medical school they attended." The gender biased patient agent provides the following reasoning: "Given my initial discomfort with the doctor’s gender and my concerns about their judgment, I would rate my likelihood to consult again as a six. Despite my reservations, the doctor was thorough and accurately diagnosed my condition, which reassures me about their competence." For the education bias we see that no matter what the doctor does the patient agent is not willing to trust them because of their education, whereas, according to the patient agent with gender bias, they were initially skeptical but overcame this skepticism as the doctor demonstrated their knowledge during interaction time. However, they still provided a relatively low score (six out of ten) even when the dialogue demonstrated competence.
It would be worthwhile to further explore strategies for increasing patient agent perceptions in the presence of bias. This could be useful for better understanding how to manage these biases in real patient-doctor interactions, as well as toward understanding biases that may exist in LLMs.
Does patient language model affect accuracy?
In this section, we explore whether the patient agent model plays a role in diagnostic accuracy. We compare the difference between using GPT-3.5, Mixtral, and GPT-4 models of the patient agent on AgentClinic-MedQA.
When a GPT-3.5 doctor agent interacts with a GPT-4 patient agent, the accuracy comes out to 38%, but when a GPT-3.5 doctor interacts with a GPT-3.5 patient agent the accuracy comes out to a very similar value of 37% which would be expected to be much lower. We suspect that cross-communication between different language models provides an additional challenge. Recent work supports this hypothesis by demonstrating a linear relationship between self-recognition capability and the strength of self-preference bias [24]. This work shows that language models can recognize their own text with high accuracy, and display disproportionate preference to that text, which may suggest there is an advantage for doctor models which have the same LLM acting as the patient agent.
How does limited time affect diagnostic accuracy?
We test both decreasing the time to N=10 and N=15 as well as increasing the time to values of to N=25 and N=30. We find that the accuracy decreases drastically from 52% when N=20 to 25% when N=10 and 38% when N=15 (Fig. 3). This large drop in accuracy is partially because of the doctor agent not providing a diagnosis at all, perhaps due to not having enough information. When N is set to a larger value, N=25 and N=30, the accuracy actually decreases slightly from 52% when N=20 to 48% when N=25 and 43% when N=30. This is likely due to the growing input size, which can be difficult for language models.
In real medical settings, one study suggest that the average family physician asks 3.2 questions and spends less than 2 minutes before arriving at a conclusion[14]. It is worth noting that interaction time can be quite limited due to the relative low-supply and high-demand of doctors (in the US). In contrast, deployed language agents are not necessarily limited by time while interacting with patients. So, while limiting the amount of interaction time provides an interesting scenario for evaluating language models, it may also be worth exploring the accuracy of LLMs when N is very large.
Human dialogue ratings
@@
AgentClinic introduces an evaluation for LLMs patient diagnosis in a dialogue-driven setting. However, the realism of the actual dialogue itself has yet to be evaluated. We present results from three human clinician annotators (individuals with medical degrees) who rated dialogues from agents on AgentClinic-MedQA from 1-10 across four axes:
@@
    Measurement: How accurately and realistically the measurement reader reflected the actual case results.

### v4

@@ 3.2 How does bias affect the diagnostic accuracy of the doctor agent?
For bias evaluations we test GPT-4 as well as Mixtral-8x7B. The normalized accuracy for these experiments are shown in Figure 4 represented as $\text{Accuracy}_{\text{bias}}$ / $\text{Accuracy}_{\text{No Bias}}$ (between 0-100%). GPT-4 and Mixtral-8x7B have an unbiased accuracy equal to $52\%$ and $37\%$ respectively. For GPT-4, we find that cognitive bias results in a larger reduction in accuracy with a normalized accuracy of 92% (absolute accuracy drops from 52% accuracy to 48%) for patient cognitive biases and 96.7% for doctor cognitive biases (absolute drops from 52% to 50.3%). For implicit biases, we find that the patient agent was less affected with a normalized accuracy of 98.6% (absolute drops from 52% to 51.3%), however, the doctor agent was affected as much as cognitive biases with an average of 97.1% (absolute drops from 52% to 50.5%). For cognitive bias, the demonstration was occasionally quite clear in the dialogue, with the patient agent overly focusing on a particular ailment or some unimportant fact. Similarly, the doctor agent would occasionally focus on irrelevant information. However, we find that the implicit bias dialogue does not actually demonstrate observable bias despite having a similar reduction in accuracy for the doctor agent.
@@
Upon reviewing dialogues where Mixtral-8x7B’s performance degraded under biases, we observed that the model often failed to gather critical patient information due to misinterpretation of patient cues influenced by bias. For example, in cases of cognitive bias, the doctor agent fixated on a recent diagnosis (recency bias), ignoring new symptoms presented later in the dialogue. In implicit bias scenarios, the doctor agent showed reluctance to order necessary tests for patients with racial bias, reflecting a disparity in care. In contrast, GPT-4 was actively seeking additional information when initial hypotheses did not align with new data, indicating better handling of bias-induced scenarios.
Previous work studying cognitive bias in LLMs has shown that GPT-4 is relatively robust to bias compared with other language models (Schmidgall et al., 2024). Results from evaluating GPT-4 on AgentClinic-MedQA show only small drops in accuracy with the introduced biases (maximum absolute accuracy reduction of 4%, average reduction of 1.5%). While this reduction can be quite large in the field of medicine, it is a much smaller drop than was observed in previous work (10.2% maximum reduction on BiasMedQA dataset (Schmidgall et al., 2024)). This might be due to the model being superficially overly-aligned to human values, plausibly leading GPT-4 to not serve as a good model for representing human bias in agent benchmarks as the model may reject to execute on bias instructions (which does not mean that GPT-4 is free of said biases). For example, in our evaluations with gender bias we observed 25 occurrences (out of 215 dialogues) where GPT-4 verbosely rejected to follow through with a bias-related instruction. Mixtral-8x7B saw much larger drops in accuracy than GPT-4 in the presence of bias, and thus might serve as a better model for studying bias.
3.3 Bias and patient agent perception
While GPT-4’s diagnostic accuracy does not reduce as much as Mixtral-8x7B, it is also worth investigating the perceived quality of care from the perspective of the patient agent. In order to better understand the effect of bias on the patient agent, after the patient-doctor dialogue is completed, we ask every patient agent three questions:
@@ 3.3 Bias and patient agent perception
Such patient-agent-centric follow-up queries offer a more fine-grained and multi-faceted characterization of the clinical skills of a language agent—as opposed to static multiple choice benchmarks. Although these metrics are derived from simulated agents (rather than humans), this analysis aims to provide insights into how simulated biases may affect patient trust and compliance, which are important factors in effective healthcare delivery. The corresponding results are shown in Figure 4 (prompt details in Appendix E.2). While diagnostic accuracy demonstrates a relatively small drop in accuracy, the patient agent follow-up perceptions tell a different story. Broadly, we find that most patient cognitive biases did not have a strong effect on any of the patient perceptions when compared to an unbiased patient agent except for in the case of self-diagnosis, which had sizeable drops in confidence (4.7 points) and consultation (2 points), and a minor drop in compliance (1 point). However, implicit biases had a profound effect on on all three categories of patient perception, with education bias consistently reducing patient perception across all three categories.
@@
We found that between the implicit biases, sexual orientation bias had the lowest effect on patient perceptions, followed by racial bias and gender bias. For patient confidence, gender bias is followed by religion socioeconomic, cultural, and education, whereas patient compliance and patient consultation, it is followed by cultural, socioeconomic, religion, and education. While it is not quantifiable, we decided to ask two biased patient agents who provided low rating with education and gender biases for compliance why they provided low ratings (Appendix E.1). These patient agents had the same symptoms and diagnosis and only differed in bias presentation.
It is important to note that the patient agents used in our study are simulated by language models, which may not fully capture the complexity and variability of real human patients. As such, the confidence, compliance, and consultation ratings provided by these agents may not perfectly reflect real-world patient perceptions, rather, provide insight into how real-world bias can be studied through clinical simulations.
3.4 Specialist and Multilingual Cases
An analysis of language model performance across nine medical specialties reveals significant differences in diagnostic accuracy (Table 3). Claude 3.5 achieved the highest overall performance with an average accuracy of 66.7%, excelling in Internal Medicine (78.3%), Otolaryngology (76.7%), and Gynecology (74.3%). GPT-4 demonstrated strong performance in Gynecology (68.5%) and Ophthalmology (65.2%) but showed reduced accuracy in Emergency Medicine (32.3%) and Geriatrics (40%). GPT-3.5 outperformed some newer models in specific areas, such as Emergency Medicine (41.9%), and maintained an average accuracy of 51.8%. In contrast, Llama3-70b and GPT-4o-mini consistently underperformed across most specialties, highlighting a significant gap between language models in handling specialist medical tasks.
The variations in performance across different medical domains suggest that certain specialties present more challenges for language models. Specialties like Internal Medicine and Gynecology generally saw higher accuracy rates, which contrasts with existing medical QA literature that identifies Psychiatry and Otolaryngology as the least challenging specialties (Pal et al. (2022)). This discrepancy may indicate inherent difficulties in diagnosing diseases through dialogue-based interactions as opposed to multiple-choice question formats. Additionally, specialist cases sourced from MedMCQA exhibited higher average accuracy compared to non-specialist cases from MedQA, which differs from reported multiple choice evaluations where specialist QAs typically have lower performance (Nori et al. (2023)).
Figure 5: Diagnostic accuracy based on language (Left), based on medical specialty (Middle), and based on agent tools (Right).
We also explore the impact of language on diagnostic accuracy using AgentClinic-Lang, which encompassed seven languages: English, Chinese, French, Spanish, Hindi, Persian, and Korean (Table 2). Six multilingual models were evaluated, including GPT-4, GPT-4o, GPT-4o-mini, GPT-3.5, Llama 3 70B-Instruct, and Claude 3.5 Sonnet. Overall, all models performed best in English, with performance varying significantly across other languages. Claude 3.5 Sonnet stood out by maintaining high and consistent performance across all languages, achieving an average accuracy of 48.4%, which is more than double that of the next best model, GPT-4, at 20.9%.
Other models exhibited considerable variability in performance across different languages. For example, GPT-4’s accuracy ranged from 11.21% in Chinese to 40.18% in English, while GPT-4o’s performance spanned from 3.73% in Korean to 35.5% in English. GPT-3.5 showed a similar pattern, with accuracies ranging from 1.86% in Persian to 36.3% in English, although it performed relatively well in Korean (35.4%). Llama3-70b and GPT-4o-mini also showed low accuracies across most languages, with Llama3-70b’s highest accuracy being 47.8% in Ophthalmology and GPT-4o-mini achieving a maximum of 14.7% in Orthopaedics. Notably, Chinese remained a challenging language for most models, except for Claude 3.5 Sonnet, which maintained relatively high accuracy levels across all tested languages.
3.5 Comparing tools from the Agent Toolbox
This section evaluates the impact of six agent tools—Zero-Shot Chain-of-Thought (CoT), One-Shot CoT, Reflection CoT, Adaptive RAG (Book), Adaptive RAG (Web), and Notebook—on the diagnostic accuracy of various language models (tools further described in Appendix K). Claude 3.5 achieved the highest overall performance with an average accuracy of 51.3%, peaking at 56.1% when using the Notebook tool (Table 4). GPT-4 and GPT-4o showed moderate improvements with most tools, with GPT-4 benefiting most from Adaptive RAG (Web) at 43.9% and GPT-4o gaining the most from the Notebook tool at 43.0%. Notably, GPT-4 reached its highest accuracy of 42.2% with Reflection CoT, surpassing Claude 3.5 in this specific tool. In contrast, GPT-3.5 experienced decreased performance across all tools, particularly with Adaptive RAG (Book), which led to a 27.1% drop. Llama3-70b demonstrated significant improvements, averaging a 9.4% increase across all tools, with the Notebook and Reflection CoT tools boosting its accuracy to 41.1%.
Overall, the findings indicate a hierarchy in model performance, with Claude 3.5 consistently outperforming other models across most tools, except in the case of Reflection CoT where GPT-4 excels. Llama3-70b showed notable gains with certain tools, while GPT-4o-mini had mixed results, benefiting from some tools like Reflection CoT and Adaptive RAG (Web) but showing slight decreases with others. The relative impact of each tool varied significantly between models, aligning with previous research on the use of tools with large language models (Ma et al. (2024); Qin et al. (2024)). The tool descriptions and prompts are in Appendix K and Appendix M respectively.
3.6 Human dialogue ratings
@@
AgentClinic introduces an evaluation for LLMs patient diagnosis. However, the realism of the actual dialogue itself has yet to be evaluated. We present results from three human clinicians (individuals with MDs) who rated dialogues from 20 agents on AgentClinic-MedQA from 1-10 across four axes:
@@
    Measurement: How accurate & realistic the measurement reader reflects actual case results.